OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

whoops sin^2x

OpenStudy (anonymous):

medal and fan for whoever helps me

OpenStudy (superdavesuper):

this one is easy - just put y=sinx and then solve y as a quadratic eqn.

OpenStudy (anonymous):

could you please show me?

OpenStudy (anonymous):

are you there?

OpenStudy (superdavesuper):

y=sinx so y^2=sin^2x....then solve for y...

OpenStudy (campbell_st):

well you have a quadratic equation. start by adding 5 to both sides of the equation and you get \[7\sin^2(x) - 14\sin(x) + 7 = 0\] now divide every term by 7 \[\sin^2(x) - 2\sin(x) + 1 = 0\] this can be factorised, its a perfect square \[(\sin(x) -1)^2 = 0\] hope this helps.... you'll need to solve it from here...

OpenStudy (anonymous):

okay that helps a lot thank you!

OpenStudy (anonymous):

:)